OpenStudy (foreverandalways13):

divide these polynomials: (2x^2-4x+2)/(x-1) PLEASE HELP

OpenStudy (foreverandalways13):

HELP

OpenStudy (loser66):

@Luigi0210

OpenStudy (loser66):

I summon him to help you because I know he can

OpenStudy (foreverandalways13):

k thanks...

OpenStudy (luigi0210):

factor the top polynomial

OpenStudy (foreverandalways13):

ummmm

OpenStudy (luigi0210):

\[\frac{ 2(x-1)^2 }{ x-1 }\]

OpenStudy (whpalmer4):

Or do long division. Observe that 2x(x-1) = 2x^2-2x, which you could subtract from the numerator, leaving 2x^2-4x+2-(2x^2-2x)= -2x+2. Now how many (x-1)'s can you get from -2x+2 (which also can be written -2(x-1))?

OpenStudy (foreverandalways13):

thanks so much both of you!!!!

OpenStudy (foreverandalways13):

u guys are awesome!!!!!

OpenStudy (luigi0210):

Naw I'm me :P